OpenStudy (anonymous):

Write the equation of the line passing through the points (-6,1) and (-4,2).

OpenStudy (anonymous):

what did we do last time? lol first you find the slope right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

1/-10

OpenStudy (anonymous):

After Finding Slope use POint Slope Form y-y1 = m (x-x1)

OpenStudy (anonymous):

(y2-y1)/(x2-x1) should have gotten you 1/2

OpenStudy (anonymous):

i think i keep using the wrong formula :( that the problem lol

OpenStudy (anonymous):

(y-y1)=((y2-y1)/(x2-x1))*(x-x1) use this formula

OpenStudy (anonymous):

just remember "y sub 2 - y sub 1 now you got your rise time time to find your run, x sub 2 - x sub 1, slope equals rise over run! lol my math teacher in 6th grade tought me that xD

OpenStudy (hba):

Let me show you a formula, \[y-y_1=\frac{ y_2-y_1 }{ x_2-x_1 }(x-x_1)\] Where, \[(x_1,y_1)=(-6,1) \ and \ (x_2,y_2)=(-4,2)\]

OpenStudy (anonymous):

did that formua confuse you? lol

OpenStudy (anonymous):

yes but i got it thank you

OpenStudy (anonymous):

alright :D yayyy good job!

OpenStudy (anonymous):

:)))))

OpenStudy (hba):

Welcome.